# Benchmarking Multi-Modal Retrieval for Long Documents

### v1

@@ 7.1. Text-Centric Document Retrieval
Current text retrieval are primarily classified into sparse and dense retrieval. For two widely-used sparse retrievers: TF-IDF (Salton et al., 1983) calculates the relevance via word frequency with the inverse document frequency, and BM25 (Robertson et al., 1994) introduces nonlinear word frequency saturation and length normalization. Dense retrievers encode content into vector representations. DPR (Karpukhin et al., 2020) is the pioneering work of dense vector representations for QA tasks. Similarly, ColBERT (Khattab and Zaharia, 2020) introduces an efficient question-document interaction model with late fine-grained term matching. Contriever (Izacard et al., 2022) leverages contrastive learning to improve content dense encoding. E5 (Wang et al., 2022) and BGE (Xiao et al., 2023) propose novel training and data preparation techniques to enhance retrieval performance. Moreover, GTE (Li et al., 2023) integrates graph-based techniques to enhance dense embedding. Most systems focus on text-based representations, neglecting the valuable visual information present in documents.
7.2. Vision-Driven Document Retrieval
@@
Vision Language Models (VLMs) (Abdin et al., 2024; Beyer et al., 2024; Bai et al., 2023; Chen et al., 2024b) have gained popularity for their ability to understand and generate text based on combined text and visual inputs. This advancement has led to the development of cutting-edge visual-driven retrievers, such as ColPali (Faysse et al., 2024) and DSE (Ma et al., 2024a). These models specifically leverage PaliGemma (Beyer et al., 2024) and Phi3-V (Abdin et al., 2024) to directly encode document page screenshots for multi-modal document retrieval. ColPali adopts similar question-document interaction as ColBERT, and represent each document page in token-level embeddings. By contrast, DSE is similar to DPR to encode each page with a single dense embedding. Visual Retrievers can directly model useful visual information, allowing for the direct utilization of multi-modal content without first converting it into text. Despite these advancements, visual retrievers face challenges, particularly in dealing with text details when document page resolutions are high. The high resolution of document pages substantially increases the computational cost and complexity of the embedding process, which may hinder the model’s performance.
7.3. Multi-modal Document Retrieval Datasets
As described in Section 1 and Table 1, there is a notable lack of a robust benchmark for multi-modal document retrieval. DocCVQA (Tito et al., 2021) is the first multi-modal document retrieval-answering task, which extracts information from a document image collection and then provides the answer. However, DocCVQA provides only 20 questions, and the corpora is limited to finance domain specifically in political disclosure. PDF-MVQA (Ding et al., 2024) is tailored for multi-modal retrieval in research articles. Note that PDF-MVQA is annotated by Chat-GPT (GPT-3.5-turbo) but not experts, and its corpora contains only biomedical articles. SciMMIR (Wu et al., 2024) proposes multi-modal retrieval in scientific research papers. However, it provides only the image-caption pairs, not the natural user query with the targeting document pages. Ma et al. (2024a) introduce two relevant datasets, namely Wiki-SS and DocMatix-IR. Wiki-SS is constructed based on Natural Questions (Kwiatkowski et al., 2019) where the evidence passage is the screenshot wikipedia webpage. Natural Questions are mainly designed for text rather than multi-modal retrieval. We also notice that the screenshot may not include the ground-truth evidence as only the front-page is screenshotted. DocMatix-IR is constructed from the largest document understanding dataset DocMatix (Laurençon et al., 2024) via filtering and hard negative mining. However, the VQA questions in DocMatix is constructed by Phi-3-small (Abdin et al., 2024) rather than experts, and these VQA questions are not de-contextualized for retrieval.
ViDoRe (Faysse et al., 2024) is the most relevant benchmark to MMDocIR. It integrates multiple DocVQA datasets (Mathew et al., 2021, 2022; Zhu et al., 2022; Li et al., 2024), and provides new documents in scientific, medical, administrative, and environment domains. Upon thorough examination of the 1,180 questions within the ViDoRe dataset, we find that more than 80% of these questions present drawbacks in terms of their complexity and context-specific nature. Typical examples include “What’s the title of the picture” and “what’s the author’s name”. They tend to be either overly simplistic or excessively dependent on a specific context which does not necessarily reflect the broader, more complex needs of effective IR. Additionally, a significant limitation of the ViDoRe dataset is its reliance on sampled document pages or cropped images rather than providing the complete document corpora.
8. Conclusion
In conclusion, multi-modal document retrieval presents a complex challenge that necessitates the integration of diverse data modalities beyond plain text. Our contributions address this challenge by introducing the MMDocIR benchmark, which includes innovative dual-task retrieval functionalities targeting both page-level and layout-level specifics of documents. The MMDocIR benchmark comprises a rich dataset featuring expertly annotated labels for 1,685 questions and bootstrapped labels for 173,843 questions, making it a pivotal resource for advancing multi-modal document retrieval for both training and evaluation purposes. Our comprehensive evaluation reveals that visual-driven retrievers significantly outperform text-driven ones, underscoring the critical importance of incorporating visual content in enhancing the effectiveness of document retrieval systems. Future work can expand on optimizing these retrieval algorithms to improve the accuracy and efficiency of multi-modal document retrieval systems.

### v2

@@ D.1 Text-Centric Document Retrieval
For text retrieval, the first step is to convert multimodal document into text using techniques, e.g., Document Parsing Chao and Fan (2004); Wang et al. (2024), Optical Character Recognition (OCR) Chaudhuri et al. (2017); Borovikov (2014); Mori et al. (1999), Layout Detection Sassioui et al. (2023); Xu et al. (2020, 2021), Chunking Chen et al. (2024a); Raina and Gales (2024); Dong et al. (2024), and Image Captioning You et al. (2016); Aneja et al. (2018). These steps are time-consuming and can introduce errors that impact the overall retrieval performance. Current text retrieval are primarily categorized as sparse or dense retrieval. For two widely-used sparse retrievers: TF-IDF Salton et al. (1983) calculates the relevance via word frequency with the inverse document frequency, and BM25 Robertson et al. (1994) introduces nonlinear word frequency saturation and length normalization. Dense retrievers encode content into vector representations. DPR Karpukhin et al. (2020) is the pioneering work of dense vector representations for QA tasks. Similarly, ColBERT Khattab and Zaharia (2020) introduces an efficient question-document interaction model with late fine-grained term matching. Contriever Izacard et al. (2022) leverages contrastive learning to improve content dense encoding. E5 Wang et al. (2022) and BGE Xiao et al. (2023) propose novel training and data preparation techniques to enhance retrieval performance. Moreover, GTE Li et al. (2023) integrates graph-based techniques to enhance dense embedding. However, most text retrieval systems overlook valuable visual information present in documents.
{tabu}
l|l|c|c|l Model Dimension Base Model HuggingFace Checkpoint
Text
DPR 768 BERT-base facebook/dpr-ctx_encoder-multiset-base facebook/dpr-question_encoder-multiset-base
[…]
ColBERT $N_{\mathrm{tok}}\times$768 BERT-base colbert-ir/colbertv2.0
[…]
Contriever 768 BERT-base facebook/contriever-msmarco
[…]
E5 1,024 BERT-large intfloat/e5-large-v2
[…]
BGE 1,024 RetroMAE BAAI/bge-large-en-v1.5
[…]
GTE 1,024 BERT-large thenlper/gte-large
Visual
DSEwiki-ss 3,072 Phi-3-Vision Tevatron/dse-phi3-v1.0
[…]
DSEdocmatix 3,072 Phi-3-Vision Tevatron/dse-phi3-docmatix-v2
[…]
ColPali $N_{\mathrm{tok}}\times$1,024 PaliGemma vidore/colpali
Table 13: Implementation details for Text and Vision Retrieval Models
D.2 Vision-Driven Document Retrieval
@@
Vision Language Models (VLMs) Abdin et al. (2024); Beyer et al. (2024); Bai et al. (2023); Chen et al. (2024b) can understand and generate text based on combined text and visual inputs. This advancement has led to the development of cutting-edge visual-driven retrievers, such as ColPali Faysse et al. (2024) and DSE Ma et al. (2024a). These models specifically leverage PaliGemma Beyer et al. (2024) and Phi3-Vision Abdin et al. (2024) to directly encode document page screenshots for multimodal document retrieval. ColPali adopts a similar question-document interaction as ColBERT, and represents each document page in token-level embeddings. By contrast, DSE is similar to DPR in that it encodes each page with a single dense embedding. Visual retrievers are capable of modeling useful visual information, allowing direct utilization of multimodal content without first converting it into text first. Despite these advancements, visual retrievers face challenges, particularly in dealing with text details when document page resolutions are high. The high resolution of document pages substantially increases the computational cost and complexity of the embedding process, which may hinder the model’s performance.
D.3 Implementation Details
In our experiments (refer to Section 5.2), we implement 9 off-the-shelf retrievers including 6 text retrievers and 3 visual retrievers. The text retrieval models deployed are namely DPR, ColBERT, Contriever, E5, BGE and GTE. These models use the WordPiece tokenizer from BERT and also inherit the maximum input length of 512 tokens from BERT Devlin et al. (2019). Additionally, we make use of the sentence-transformer library [^12] https://www.sbert.net/ when deploying E5, BGE and GTE. The visual retrieval models deployed are namely DSEwiki-ss, DSEdocmatix, and ColPali. We use pre-trained checkpoints available on HuggingFace  [^13] https://huggingface.co/; the specific checkpoint information can be found in Table D.1 alongside other configuration details.
Appendix E Dataset Demonstration
E.1 Document Pages by Domains
The documents in MMDocIR can be categorized into 10 types. We provide examples of each type as below.
  • Admin & Industry: These documents primarily consist of instructional and overview content on industry, reflected by the dominance of text-based questions (78.0%) and a smaller reliance on visual evidence (image questions only 20.3%), which shows a text-heavy structure (70.1%). Some detailed examples are shown in Figure 5(b).
  • Tut & Workshop: Documents in this category focus on slides or tutorials, which exhibit a balanced question modality: 61.7% text, 24.5% image, and 9.5% table questions. Strong visual components are present, with 57.4% of its content being images—the highest among all categories. Some detailed examples are shown in Figure 5(c).
  • Academic Paper: These documents are formal publications with structured layouts, citations, and academic pictures. The questions span multiple modalities: 28.8% text, 25.7% image, and 50.0% table. Text modality dominates content distribution (74.6%), with the presence of tables (11.1%) and images (12.8%) demonstrating rich multimodal alignment and explicit questions with answers. Some detailed examples are shown in Figure 6(a).
  • Brochure: Designed for promotional purposes, the brochure category contains highly visual documents. Over 52.6% of questions are image-based—the highest among all domains—while text-based questions account for only 60.5%. Modality distribution is similarly diverse: 50.8% image, showcasing their visually complex layout. Some detailed examples are shown in Figure 6(b).
  • Financial Report: These documents involve massive numerical and quantitative data, reflected in a high proportion of table questions (54.5%) and strong table content distribution (29.2%). While text remains significant (60.3%), the inclusion of tabular and numerical analysis is essential for understanding these documents. Some detailed examples are shown in Figure 6(c).
  • Guidebook: Instruction manuals for electronics and tools, guidebooks exhibit the most balanced question modality: 51.8% text, 54.4% image, and 26.8% table, indicating multimodal instructional designs. Some detailed examples are shown in Figure 7(a).
  • Government: This category covers policy files and governmental reports. It is highly text-centric with 69.9% text questions and 88.2% text content. This reflects the formal and regulatory nature of such documents. Some detailed examples are shown in Figure 7(b).
  • Laws: Legal documents exhibit strong textual dominance both in questions (62.1%) and content (83.8%), with very limited visual presence (image content only 1.6%). They often maintain specific formats and focus on linguistic interpretation rather than visual layout. Some detailed examples are shown in Figure 7(c).
  • News: Although based on only one document, the “News” domain shows notable multimodal richness. It includes a significant image portion (39.8%), high text presence (48.5%), and 11.6% titles. This reflects the use of images and headlines typical of news articles. Some detailed examples are shown in Figure 7(d).
E.2 Document Layouts
In this section, we present 9 pages along with their detected layouts, which are highlighted for better visualizations, as shown in Figure 8, 9, and 10. Specifically, layout detection identifies the spatial location of different content types, such as images, tables, and text within a document. With the help of layout detection, we can precisely locate the specific position of an answer, whether it is an image, a text paragraph, or a table. This enables a more fine-grained layout-level evaluation of multimodal retrieval capabilities.
E.3 Annotation Examples
In this section, we present 4 annotation examples that illustrate typical multimodal retrieval and reasoning patterns, which help explain the construction and retrieval process. Each annotation includes the following primary components: question, answer, page-level labels, and layout-level labels. The page-level labels show the selected pages that contain ground truth evidence. Based on these selected pages, layout-level labels further display the specific layout box detection of evidence. These examples frequently require reasoning across multiple pages and modalities. The evidence encompasses diverse formats such as figures, charts, tables, and texts, highlighting the complexity and richness of the multimodal retrieval tasks.
(a) Page screenshots in Research Report domain.
(b) Page screenshots in Administration & Industry domain.
(c) Page screenshots in Tutorial & Workshop domain.
Figure 5: The screenshot examples of typical document pages for (a) Research Report, (b) Administration & Industry, and (c) Tutorial & Workshop domain.
(a) Page screenshots in Academic Paper domain.
(b) Page screenshots in Brochure domain.
(c) Page screenshots in Financial Report domain.
Figure 6: The screenshot examples of typical document pages for (a) Academic Paper, (b) Brochure, and (c) Financial Report domain.
(a) Page screenshots in Guidebook domain.
(b) Page screenshots in Government domain.
(c) Page screenshots in Laws domain.
(d) Page screenshots in News domain.
Figure 7: The screenshot examples of typical document pages for (a) Guidebook, (b) Government, (c) Laws domain, and (d) News domain.
(a) Example 1: original page vs. page highlighted with layout bounding boxes.
(b) Example 2: original page vs. page highlighted with layout bounding boxes.
(c) Example 3: original page vs. page highlighted with layout bounding boxes.
Figure 8: The 3 examples illustrate the function and effectiveness of layout detection on document pages.
(a) Example 1: original page vs. page highlighted with layout bounding boxes.
(b) Example 2: original page vs. page highlighted with layout bounding boxes.
(c) Example 3: original page vs. page highlighted with layout bounding boxes.
Figure 9: The 3 examples illustrate the function and effectiveness of layout detection on document pages.
(a) Example 1: original page vs. page highlighted with layout bounding boxes.
(b) Example 2: original page vs. page highlighted with layout bounding boxes.
(c) Example 3: original page vs. page highlighted with layout bounding boxes.
Figure 10: The 3 examples illustrate the function and effectiveness of layout detection on document pages.
Figure 11: This example shows a typical image retrieval and reasoning task that requires synthesizing information from pie chart.
Figure 12: This example shows a typical multi-page retrieval task that requires synthesizing information from text passages across multiple pages.
Figure 13: This example shows a typical image reasoning task that requires synthesizing information from specific image.
Figure 14: This example shows a typical multi-page image and text reasoning task that requires synthesizing cross-modal information from image and text.
Sub dataset #Not #Barely #Suit- #Total HuggingFace Resource
suitable suitable able
arXiVQA 245 203 52 500 vidore/arxivqa_test_subsampled
DocVQA 345 130 25 500 vidore/docvqa_test_subsampled
InfoVQA 139 284 77 500 vidore/infovqa_test_subsampled
TAT-DQA 373 121 6 500 vidore/tatdqa_test
Industrial 78 260 62 400
vidore/syntheticDocQA_energy_test
vidore/syntheticDocQA_healthcare_industry_test
vidore/syntheticDocQA_government_reports_test
vidore/syntheticDocQA_artificial_intelligence_test
Sum 1180 998 222 2,400 -
Percentage 49.1% 41.5% 9.25% - -
Table 14: Document statistics for ViDoRe Benchmark.
Appendix F Detailed Analysis of ViDoRe Benchmark
F.1 Query and Annotation Analysis
As mentioned in Section 6, ViDoRe Faysse et al. (2024) is the most relevant benchmark to MMDocIR. It integrates several datasets such as DocVQA Mathew et al. (2021), InfoVQA Mathew et al. (2022), TAT-DQA Zhu et al. (2022), arXiVQA Li et al. (2024), and providing new documents in scientific, medical, administrative, and environment domains. In this Appendix, we elaborate our analysis of the sampled 2,400 questions sampled from ViDoRe. The statistics are shown in Table 14. ViDoRe test set contains questions in either English or French. In our work, we examine only the English questions. For academic subsets, we examine all 1,500 questions: 500 questions from DocVQA, 500 questions from InfoVQA, and 500 questions from arXiVQA. In TAT-DQA, we sample and examine the first 500 questions. For the industrial documents, we select 100 questions from each domain (i.e., energy, healthcare, government, and artificial intelligence). We examine sampled questions and summarize these questions into 3 categories:
  • Unsuitable Queries. Queries that are not well-suited for IR systems can often burden these systems by generating numerous irrelevant results. For example, a query such as “What’s the x-axis of the figure” is likely to prompt matches from multiple passages within document corpora that mention figures with an x-axis. This tends to happen because the query is overly broad and lacks contextual specificity. When such queries stem from Document Visual Question Answering (DocVQA) tasks targeting a single image, the challenge is exacerbated, as the reliance on precise context increases while the target remains too vague, undermining the fundamental principles of effective IR.
  • Barely Suitable Queries. Queries that fall into this category provide some guidance towards locating useful passages, yet suffer from a lack of precise detail. These queries often fetch moderate number of passages, where both relevance and focus may not be as sharp. For example, the query “What was the total assets from AMER in 2018?” is meant for Visual Question Answering (VQA) focused on a specific financial topic. Although this seems specific, the issue arises when multiple sections within an annual report discuss AMER’s total assets for the year. This causes significant confusion since ViDoRe is set to acknowledge only a single passage as the verified answer. This lack of uniqueness in the ground truth makes it hard to evaluate the actual performance of IR system.
  • Suitable Queries. The most effective queries for IR systems are characterized by their specificity and ability to distinguish between different sections of texts. These queries often involve precise facts or detailed inquiries that facilitate pinpointing exact passages. For instance, the question “What was the magnitude of the earthquake that occurred in Maule on 2/27/2010?” incorporates significant keywords and details that guide the retrieval system directly to the necessary data. Such queries align perfectly with the objectives of IR, leveraging specificity and detailed context to efficiently retrieve most relevant information.
The comprehensive analysis of our queries, as presented in Table 14, reveals a significant challenge in adapting questions from Visual Question Answering (VQA) datasets (such as DocVQA, InfoVQA, TAT-DQA, and arXiVQA) for Information Retrieval (IR) purposes. Only 8% of these queries prove suitable for effective IR usage. In comparison, queries derived from industrial documents perform slightly better, with 15.5% deemed suitable. A common issue identified is that these queries are either excessively simplistic or highly specific to a particular context. Our findings suggest that the primary difficulty stems from the inherent differences between DocIR and DocVQA. VQA queries are typically crafted to address content on a specific page or within a particular image, inherently limiting their scope and specificity. This specificity and simplism are functional within the confines of the intended VQA context but pose substantial limitations when such queries are repurposed for IR tasks. Due to this gap, we exclude ViDoRe benchmark from our experiments.
F.2 Document Corpora Analysis
ViDoRe bootstrap document corpora directly from existing DocVQA benchmarks (i.e., DocVQA, InfoVQA, TAT-DQA, arXiVQA) that perform single-page VQA. In the DocVQA setting, only selected pages are provided for VQA, rather than the entire document pages. For arXivQA, the retrieved passages are not document pages, but are cropped images (e.g., figures, tables, and charts). In our experiments, we need to rely on the entire documents pages to evaluate retrieval on long documents. To bridge the gap of missing complete document corpora, we put in considerable efforts to collect the original documents of existing DocVQA datasets, as mentioned in Section 4.1.
Appendix G License Agreements
We ensure that the distribution of each dataset complies with the corresponding licenses, all of which are listed below:
  • MMLongBench-Doc: is under Apache-2.0 license agreement for academic research purposes.
  • DocBench: we achieved the agreement of usage as academic research from the dataset’s author.
  • MP-DocVQA: is under “MIT License” license agreement for academic research purposes.
  • SlideVQA: is under “NTT License” license agreement for academic research purposes.
  • TAT-DQA: is under “CC-BY-4.0” license agreement for academic research purposes.
  • ArXivQA: is under “CC-BY-SA-4.0” license agreement for academic research purposes.
  • SciQAG: is under “CC-BY-4.0” license agreement for academic research purposes.
  • DUDE: is under “GPL-3.0” license agreement for academic research purposes.
  • CUAD: is under “CC-BY-4.0” license agreement for academic research purposes.
For the new annotations contributed in MMDocIR, including but not limited to the questions, page and layout annotations, we make them available solely for research purposes. Users are permitted to use, modify, and share these annotations for academic and non-commercial research activities. Any other use, including commercial exploitation, is not permitted without explicit written permission from the authors.
Appendix H Ethical Considerations
The introduction and broader adoption of MMDocIR may have potential ethical impacts spanning both positive and negative dimensions. Below, we outline possible negative consequences and discuss potential mitigation strategies:
Privacy Risks: MMDocIR enables models to retrieve relevant information over lengthy, multimodal documents, which may include sensitive personal, financial, or health information. There is a risk that such technologies could be leveraged for large-scale surveillance, unauthorized extraction of personal data, or other privacy violations.
Fairness and Bias: If benchmarked models are trained or evaluated on data that does not reflect diverse demographic, linguistic, and backgrounds, outputs may exhibit biases. This may lead to unfair decision-making or stereotypes.
Mitigation Strategies: To mitigate these risks, we make sure that: (i) Benchmark development uses only publicly available, carefully vetted datasets, with sensitive information anonymized or removed; (ii) Retrieval outputs are monitored for bias and fairness.
We encourage researchers and practitioners employing MMDocIR to be mindful of these factors and to actively work toward responsible development and deployment, including transparency about limitations and proactive safeguards where needed. We welcome community feedback and collaboration on best practices to further reduce risks as this technology evolves.